• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Holmstrom

    Log entry time 16:04:52 on December 3,2009

    Entry number 304409

    keyword=Day Shift Summary

    8:00 Shift starts taking production data. LHRS DAQ is down. Experts working on it.


    8:10 LHRS DAQ fixed.


    10:01 Request Access for angle change.


    12:54 Request 2uA raster OFF
    12:56 Request a HARP scan
    13:10 HARP B looks good. Fit for HARP B fails.
    13:14 Moved to BeO. Ask for 5uA. OK
    13:16 Ask for "4x4" raster and 10uA.
    13:22 Spot++ looks good.
    13:22 Turn on VDCs.
    13:23 Move target to LD2
    13:35 Request 2uA on LD2
    13:38 Checked rates and Run Spot++ looks good.
    13:40 ask for 6x6 raster and check spot++
    13:42 Start HRS runs for Q2 measurement.
    13:47 Spot ++ looks good
    13:51 Finished the Q2 measurement.
    13:51 Turn off VDCs to go to production.
    13:59 VDCs off ramping up to production.

    14:03 Start Production




    Parity Runs
    14386 7:55-8:52 LD2 100k
    14387 8:53-9:50 LD2 100k
    14388 9:51-10:03 LD2 23k
    14389 14:03-14:59 LD2 100k
    14390 15:01-


    LHRS Runs
    25947 Junk
    25948 8:53-9:50
    25949 9:51-10:03

    25952 Q2 run 13:41-13:50 700k

    25954 14:03-14:59
    25955 15:01-


    RHRS Runs

    5480 7:55-8:52
    5481 8:53-9:50
    5482 9:51-10:03

    5484 Q2 run 13:43-13:50 500k

    5486 14:03-14:59
    5487 15:01-